Molly consumes two goods, good x and good y and her preferences are represented by the utility function U (x, y) = 1/2x^2 + 4y.

1. Draw (sketch) Molly’s indifference curves for U(x,y) = 10, U(x,y) = 16, U(x,y) = 24 and for U(x,y) = 32.5.

2. Do Molly’s preferences satisfy strict monotonicity? Explain briefly

3. Do the indifference curves you’ve drawn reflect preferences that are convex? Explain briefly
Answer #1

(n,y)= lu+4y. This and is the preferences case of can be cuasi- linear preferences represented as 14=32-5. 19:२५ 719=16 igelo